Could you please help me with this problem: The ages of three brothers are consecutive even integers. Three times the age of the youngest brother exceeds the oldest brother's age by 48 years. What is the age of the youngest brother?
Answer by rwm(914)

x;x+1;x+2
3*x=x+2+48
2x=50
x=25
check
3*25=27+48
75=75
ok